The following data pertain to Aurora Electronics for the month of February.

Static Budget Actual
Units sold 14,000 11,000
Sales revenue $ 196,000 $ 137,500
Variable manufacturing cost 5,000 42,000
Fixed manufacturing cost 25,000 25,000
Variable selling and administrative cost 15,000 9,000
Fixed selling and administrative cost 15,000 15,000

  

Required:

Compute the sales-price and sales-volume variances for February. (Indicate the effect of each variance by selecting "Favorable" or "Unfavorable". Select "None" and enter "0" for no effect (i.e., zero variance).)


Sales-price variance Sales-volume variance
Answer #1

Sales price Variance = AS-(SP*AQ)

= 137,500 - (11,000*(196,000/14,000))

= 16,500 Unfavorable

Sales volume variance = (SQ-AQ) *SP

= (14,000 - 11,000) * (196,000/14,000)

= 42,000 Unfavorable